OpenStudy (anonymous):

Find the integral of x^5 cos (x^3) dx

OpenStudy (cggurumanjunath):

use integration by parts !

OpenStudy (cggurumanjunath):

ilate rule !

OpenStudy (anonymous):

I know, but I'm having troule on the antiderivative of v

OpenStudy (agent0smith):

x^3*(x^2cos(x^3)) and integrate by parts

OpenStudy (cggurumanjunath):

u=x^5 v=cos(x^3)

OpenStudy (agent0smith):

How will that v work?

OpenStudy (cggurumanjunath):

d/dx(sinx)=cosx

OpenStudy (cggurumanjunath):

antiderivative of cos x =sinx

OpenStudy (cggurumanjunath):

antiderivative =integral.

OpenStudy (anonymous):

for my u i picked x^3 and for my du= 3x^2

OpenStudy (cggurumanjunath):

u=x^5 !

OpenStudy (anonymous):

why pick u= x^5?

OpenStudy (cggurumanjunath):

because we haVE TO FIND the integral of x^5 cos (x^3) .

OpenStudy (cggurumanjunath):

u=x^5 v=cos(x^3)

zepdrix (zepdrix):

u=x^3, dv=x^2 cos x^3 dx I think you were on the right track. Not sure what guru is talking about :o

OpenStudy (agent0smith):

^yep the reason you do it that way is because you can easily use u-sub on x^2 cos(x^3)

OpenStudy (agent0smith):

You might have to integrate it by parts more than once though, i haven't worked through it yet, but sometimes in these cases you need to.

OpenStudy (anonymous):

how do I find v?

zepdrix (zepdrix):

You'll need to make a substitution within your v, it's a little tricky. Let m=x^3 or something like that. ( We don't want to use u again).

OpenStudy (anonymous):

can you draw that? I'm a visual learner.

zepdrix (zepdrix):

v = int x^2 cos x^3 dx u = x^3 ----> (1/3)du = x^2 dx v = int cos u (1/3)du Unfortunately the drawing tool isn't working right now :( Otherwise I would..

OpenStudy (anonymous):

ok, I'm still confused but thanks for your help!

zepdrix (zepdrix):

I did a sneaky little trick there. I multiplied by 3 and left a factor of 1/3 outside of the integral. Our dv will change to v a little easier that way. I may have skipped a few steps in there :o lemme know if it's not clear.

OpenStudy (anonymous):

where did the 1/3 come from?

OpenStudy (loser66):

to me, I let u = x^6 --> x^3 = sqrt (u) and du = 6x^5 dx so, the integral = 1/6 integral cos (sqrt u) du

OpenStudy (agent0smith):

^ that might be just as awkward to integrate

OpenStudy (loser66):

hehehe... may be!! http://www.wolframalpha.com/input/?i=integral+%28cos+%28sqrt+%28u%29%29%29du

OpenStudy (agent0smith):

http://www3.wolframalpha.com/Calculate/MSP/MSP7051ga2a41igd48409400003ii2a968aed0fh3h?MSPStoreType=image/png&s=45&w=458&h=648 maybe it's easier than the other method, but i haven't tried the other yet